stylequeen:

19. (03.06 LC) Choose the equation below that represents the line passing through the point (−5, 1) with a slope of three halves. (1 point) y − 5 = three halves (x + 1) y + 1 = three halves (x − 5) y + 5 = three halves (x − 1) y − 1 = three halves (x + 5)

darkknight:

point slope form (-5,1) and slope = 3/2 \[y-1=3/2(x+5)\] That is how it will look like

stylequeen:

ok thank u soo much

stylequeen:

4

darkknight:

yep np, have a good day

stylequeen:

u2